# Sample-Efficient Reinforcement Learning from Human Feedback via Information-Directed Sampling

### v1

@@ A-D Proof of Proposition 1
where the first inequality is due to triangle inequality; the second inequality again uses triangle inequality, and the fact that $P_{h}^{\mathcal{E}}(x|o)\leq 1/\beta$; the third inequality is due to the fact that $|a-b|\leq B\cdot|\log\frac{a}{b}|$ for any $a,b\in(0,B)$ and $\bigg{|}\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(\cdot|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|o)}\bigg{|}\leq 2\log\frac{1}{\beta}$. Note that when $P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|o)$ equals zero, since ¯$\bar{\mathcal{E}}_{t}$ is the mean MDP, it turns out that $P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(\cdot|o)$ also equals zero, yielding $\bigg{|}\log\frac{P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(\cdot|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|o)}\bigg{|}=0.$ Finally, the last inequality is due to Eqn. (11).
For the second term in Eqn. (A-D), adopting a similar approach, we have
$\bigg{|}\int_{[0,1]}R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)\log\frac{R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-R_{h}^{\mathcal{E}}(x|o)\log\frac{R_{h}^{\mathcal{E}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
$\leq\bigg{|}\int_{[0,1]}R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)\log\frac{R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-R_{h}^{\mathcal{E}}(x|o)\log\frac{R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}+\bigg{|}\int_{[0,1]}R_{h}^{\mathcal{E}}(x|o)\log\frac{R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-R_{h}^{\mathcal{E}}(x|o)\log\frac{R_{h}^{\mathcal{E}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
@@
$\leq\int_{[0,1]}\left|R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)-R_{h}^{\mathcal{E}}(x|o)\right|\cdot\bigg{|}\log\frac{R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\bigg{|}\mathrm{d}x+\int_{[0,1]}B\cdot\bigg{|}\log\frac{R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x$
@@
$\leq B\cdot\left(1+2\log\frac{1}{\beta}\right)\int_{[0,1]}\bigg{|}\log\frac{R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x$
$\leq\frac{(1-2\log{\beta})\epsilon}{2H}.$ (57)
Hence, adding up Eqn. (A-D) and Eqn. (A-D), we obtain
$|r^{\prime}_{h}(s_{h},a_{h})-\bar{r}_{h}(s_{h},a_{h})|\leq\frac{\lambda}{2}\cdot\frac{(1-2\log{\beta})\epsilon}{H}.$
@@
$\bigg{|}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})\bigg{]}-\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\bigg{|}\leq\frac{\lambda}{2}\cdot\epsilon\cdot(1-2\log\beta)$ (58)
@@ Appendix B Basic Properties of the Measure $\ell_{g}$
The following result deals with the problem of convexity. Let $B(\mathcal{C},\epsilon)$ be an $\epsilon$-ball with its center at $\mathcal{C}$. Note that $B(\mathcal{C},\epsilon)$ is not essentially convex under $\ell_{g}$: for $P,Q\in B(\mathcal{C},\epsilon)$ and $\lambda\in(0,1)$, it does not hold that $\lambda P+(1-\lambda)Q\in B(\mathcal{C},\epsilon)$. However, using Lemma 1, we have the following result:
Lemma 3.
For any $P,Q\in B(\mathcal{C},\epsilon)$ and $\lambda\in[0,1]$, $\lambda P+(1-\lambda)Q\in B(\mathcal{C},\epsilon)$ lies in the $(2\epsilon)$-ball at $\mathcal{C}$, i.e.,
@@
$\ell_{g}(\lambda P+(1-\lambda)Q,\mathcal{O})$ $=\sup_{o}\int_{x}\bigg{|}\log\frac{\lambda P(x|o)+(1-\lambda)Q(x|o)}{\mathcal{C}(x|o)}\bigg{|}$
@@
where the first inequality uses the fact that $|\log(\lambda a+(1-\lambda)b)|\leq|\log a|+|\log b|$ for any $a,b>0$ and $\lambda\in[0,1]$; the second inequality is due to $P,Q\in B(\mathcal{C},\epsilon)$. This finishes the proof of Lemma 3. ∎
@@ Appendix C Technical Lemmas
C-A Value Function and Mutual Information
Lemma 4.
@@ C-A Value Function and Mutual Information
$V_{1,\pi}^{\mathcal{E}}(s_{1})-V_{1,\pi}^{\mathcal{E}^{\prime}}(s_{1})=$ $\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\mathcal{E}}\otimes R_{h}^{\mathcal{E}})(\cdot|s_{h},a_{h})}\left[r^{\prime}+V_{h+1,\pi}^{\mathcal{E}}(s^{\prime})\right]$ (59)
@@
Lemma 5.
The mutual information $\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$ can be lower bounded as
$\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\geq\sum_{h=1}^{H}\mathbb{E}_{t}\left[\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\left[D_{\text{KL}}\left((P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\tilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)\right]\right].$ (60)
Proof of Lemma 5.
@@
$\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$ $=\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(s_{h}^{t,1},a_{h}^{t,1},r_{h}^{t,1},s_{h}^{t,0},a_{h}^{t,0},r_{h}^{t,0})\mid(\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1})\right)+\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};o_{t}\mid(\mathcal{H}_{t,H},\mathcal{R}_{t,H})\right)$ (61)
$=\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\tilde{\mathcal{E}}_{t}^{*};s_{h}^{t,1}\mid\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\tilde{\mathcal{E}}_{t}^{*};a_{h}^{t,1}\mid s_{h}^{t,1},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}$
$\quad+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\tilde{\mathcal{E}}_{t}^{*};r_{h}^{t,1}\mid s_{h}^{t,1},s_{h}^{t,1},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\tilde{\mathcal{E}}_{t}^{*};s_{h}^{t,0}\mid s_{h}^{t,1},a_{h}^{t,1},r_{h}^{t,1},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}$
$\quad+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\tilde{\mathcal{E}}_{t}^{*};a_{h}^{t,0}\mid s_{h}^{t,1},a_{h}^{t,1},r_{h}^{t,1},s_{h}^{t,0},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}$
$\quad+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\tilde{\mathcal{E}}_{t}^{*};r_{h}^{t,0}\mid s_{h}^{t,1},a_{h}^{t,1},r_{h}^{t,1},s_{h}^{t,0},a_{h}^{t,0},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}+\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};o_{t}\mid(\mathcal{H}_{t,H},\mathcal{R}_{t,H})\right).$
From  [7], the first three terms on the right side of the above equation are equal to
$\sum_{h=1}^{H}\mathbb{E}_{t}\left[\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\left[D_{\text{KL}}\left((P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}\otimes r_{h}^{\tilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes r_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)\right]\right].$ (62)
@@
C-B Posterior Consistency
Lemma 6.

### v3

@@ Appendix A Proofs of Theorem and Proposition
where the first inequality is due to triangle inequality; the second inequality again uses triangle inequality, and the fact that $P_{h}^{\mathcal{E}}(x|o)\leq B$; the third inequality is due to the fact that $|a-b|\leq B\cdot|\log\frac{a}{b}|$ for any $a,b\in(0,B)$ and $\bigg{|}\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(\cdot|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|o)}\bigg{|}\leq\log\frac{B}{\beta}$. Note that when $P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|o)$ equals zero, since $\bar{\mathcal{E}}_{t}$ is the mean MDP, it turns out that $P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(\cdot|o)$ also equals zero, yielding $\bigg{|}\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(\cdot|o)}{P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|o)}\bigg{|}=0.$ Finally, the last inequality is due to Eqn. (4.10).
If we employed the KL divergence or $\ell_{1}$ metric, the derivation of third inequality would become invalid.
For the second term in Eqn. (A.4), adopting a similar approach, we have
$\bigg{|}\int_{[0,1]}R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)\log\frac{R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-R_{h}^{\mathcal{E}}(x|o)\log\frac{R_{h}^{\mathcal{E}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
$\leq\bigg{|}\int_{[0,1]}R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)\log\frac{R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-R_{h}^{\mathcal{E}}(x|o)\log\frac{R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
$\qquad+\bigg{|}\int_{[0,1]}R_{h}^{\mathcal{E}}(x|o)\log\frac{R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}-R_{h}^{\mathcal{E}}(x|o)\log\frac{R_{h}^{\mathcal{E}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\mathrm{d}x\bigg{|}$
@@
$\leq\int_{[0,1]}\left|R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)-R_{h}^{\mathcal{E}}(x|o)\right|\cdot\bigg{|}\log\frac{R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\bar{\mathcal{E}}_{t}}(x|o)}\bigg{|}\mathrm{d}x+\int_{[0,1]}B\cdot\bigg{|}\log\frac{R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x$
@@
$\leq B\cdot\left(1+\log\frac{B}{\beta}\right)\int_{[0,1]}\bigg{|}\log\frac{R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{R_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x$
$\leq\frac{(1+\log(B/\beta))\epsilon}{2H}.$ (A.32)
Hence, adding up Eqn. (A.4) and Eqn. (A.4), we obtain
$|r^{\prime}_{h}(s_{h},a_{h})-\bar{r}_{h}(s_{h},a_{h})|\leq\frac{\lambda}{2}\cdot\frac{(1+\log(B/\beta))\epsilon}{H}.$
@@
$\bigg{|}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})\bigg{]}-\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\bigg{|}\leq\frac{\lambda}{2}\cdot\epsilon\cdot(1+\log\frac{B}{\beta})$ (A.33)
@@
A.5 Proof of Proposition 5.4
The proof follows essentially the same structure as that of Proposition 1, with the only difference lying in the third inequality of Eq. (A.4). By applying the Mean Value Theorem, we have
$\int_{\mathcal{S}}\bigg{|}\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x\leq\frac{1}{\beta}\int_{\mathcal{S}}\bigg{|}P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)-P_{h}^{\mathcal{E}}(x|o)\bigg{|}\mathrm{d}x=\frac{1}{\beta}\ell_{1}(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{E}})\leq\frac{1}{\beta}\cdot\text{constant}\cdot\epsilon.$
We can derive the third inequality in Eq. (A.4) as
$B\cdot\left(1+\frac{B}{\beta}\right)\int_{\mathcal{S}}\bigg{|}\log\frac{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|o)}{P_{h}^{\mathcal{E}}(x|o)}\bigg{|}\mathrm{d}x\leq\frac{1}{\beta}\cdot\text{constant}\cdot\epsilon.$
In the conclusion of Proposition 1, the term involving $\beta$ has been modified from $\log\frac{1}{\beta}$ to $\frac{1}{\beta}$. Therefore, with a suitably chosen partition radius, we have
$\bigg{|}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})\bigg{]}-\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\bigg{|}\leq\frac{\lambda}{2}\cdot\epsilon\cdot(1+\frac{B}{\beta})$ (A.34)
@@ Appendix B Basic Properties of the Measure $\ell_{g}$
The following result deals with the problem of convexity. Let $B(\mathcal{C},\epsilon)$ be an $\epsilon$-ball with its center at $\mathcal{C}$. Note that $B(\mathcal{C},\epsilon)$ is not essentially convex under $\ell_{g}$: for $P,Q\in B(\mathcal{C},\epsilon)$ and $\lambda\in(0,1)$, it does not hold that $\lambda P+(1-\lambda)Q\in B(\mathcal{C},\epsilon)$. However, using Lemma 4.3, we have the following result:
Lemma B.1.
For any $P,Q\in B(\mathcal{C},\epsilon)$ and $\lambda\in[0,1]$, $\lambda P+(1-\lambda)Q$ lies in the $(2\epsilon)$-ball at $\mathcal{C}$, i.e.,
@@
$\ell_{g}(\lambda P+(1-\lambda)Q,\mathcal{C})$ $=\sup_{o}\int_{x}\bigg{|}\log\frac{\lambda P(x|o)+(1-\lambda)Q(x|o)}{\mathcal{C}(x|o)}\bigg{|}$
@@
where the first inequality uses the fact that $|\log(\lambda a+(1-\lambda)b)|\leq|\log a|+|\log b|$ for any $a,b>0$ and $\lambda\in[0,1]$; the second inequality is due to $P,Q\in B(\mathcal{C},\epsilon)$. This finishes the proof of Lemma B.1. ∎
B.1 A Counterexample to Demonstrate Non-convex
We present the following counterexample to demonstrate that the unit ball under $\ell_{g}$ is not convex. Let $X$ be the set of positive and Lebesgue integrable functions defined on $[0,1]$. We show that the ball centered at $f_{0}\in X$
$B(f_{0},r)=\bigl{\{}f\in X:\,\ell_{g}(f,f_{0})\leq r\bigr{\}}$
is not convex. Let
$f(x)=\begin{cases}5&0\leq x\leq 0.1,\\ \frac{5}{9}&0.1<x\leq 1,\end{cases}\qquad g(x)=\begin{cases}\frac{5}{9}&0\leq x\leq 0.9,\\ 5&0.9<x\leq 1,\end{cases}\qquad f_{0}(x)=1.$
It can be verified:
$\int_{0}^{1}f(x)\,dx=0.1\cdot 5+0.9\cdot\frac{5}{9}=1,$
and the same holds for $g$. Hence $f,g$ can be seen as the two probability measures. We also have
$\ell_{g}(f,f_{0})=0.1\log 5+0.9\log\!\Bigl{(}\tfrac{9}{5}\Bigr{)}\approx 0.68995.$
The same holds for $\ell_{g}(g,f_{0})$. And
$\ell_{g}(\frac{f+g}{2},f_{0})=\ell_{g}(\frac{25}{9},1)=|\log\frac{25}{9}|\approx 1.02165.$
Let $r=1$, then $\ell_{g}(f,f_{0})=\ell_{g}(g,f_{0})\approx 0.68995<r,\ell_{g}(\frac{f+g}{2},f_{0})\approx 1.02165>r$.
So, $f,g\in B(f_{0},r)$ but $\frac{f+g}{2}\notin B(f_{0},r)$.
@@ Appendix C Technical Lemmas
C.1 Value Function and Mutual Information
We cite the following lemma from  Moradipari et al. (2023). Similar results can be found in  Osband et al. (2013); Foster et al. (2021).
Lemma C.1.
@@ C.1 Value Function and Mutual Information
$V_{1,\pi}^{\mathcal{E}}(s_{1})-V_{1,\pi}^{\mathcal{E}^{\prime}}(s_{1})=$ $\sum_{h=1}^{H}\mathbb{E}_{\pi}^{\mathcal{E}^{\prime}}\bigg{[}\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\mathcal{E}}\otimes R_{h}^{\mathcal{E}})(\cdot|s_{h},a_{h})}\left[r^{\prime}+V_{h+1,\pi}^{\mathcal{E}}(s^{\prime})\right]$ (C.1)
@@
Lemma C.2.
The mutual information $\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$ can be lower bounded as
$\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\geq\sum_{h=1}^{H}\mathbb{E}_{t}\left[\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\left[D_{\text{KL}}\left((P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)\right]\right].$ (C.2)
Proof of Lemma C.2.
@@
$\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$ (C.3)
$=\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(s_{h}^{t,1},a_{h}^{t,1},r_{h}^{t,1},s_{h}^{t,0},a_{h}^{t,0},r_{h}^{t,0})\mid(\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1})\right)+\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};o_{t}\mid(\mathcal{H}_{t,H},\mathcal{R}_{t,H})\right)$
$=\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\widetilde{\mathcal{E}}_{t}^{*};s_{h}^{t,1}\mid\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\widetilde{\mathcal{E}}_{t}^{*};a_{h}^{t,1}\mid s_{h}^{t,1},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}$
$\quad+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\widetilde{\mathcal{E}}_{t}^{*};r_{h}^{t,1}\mid s_{h}^{t,1},s_{h}^{t,1},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\widetilde{\mathcal{E}}_{t}^{*};s_{h}^{t,0}\mid s_{h}^{t,1},a_{h}^{t,1},r_{h}^{t,1},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}$
$\quad+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\widetilde{\mathcal{E}}_{t}^{*};a_{h}^{t,0}\mid s_{h}^{t,1},a_{h}^{t,1},r_{h}^{t,1},s_{h}^{t,0},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}$
$\quad+\sum_{h=1}^{H}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\big{(}\widetilde{\mathcal{E}}_{t}^{*};r_{h}^{t,0}\mid s_{h}^{t,1},a_{h}^{t,1},r_{h}^{t,1},s_{h}^{t,0},a_{h}^{t,0},\mathcal{H}_{t,h-1},\mathcal{R}_{t,h-1}\big{)}+\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};o_{t}\mid(\mathcal{H}_{t,H},\mathcal{R}_{t,H})\right).$
From  Moradipari et al. (2023), the first three terms on the right side of the above equation are equal to
$\sum_{h=1}^{H}\mathbb{E}_{t}\left[\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\left[D_{\text{KL}}\left((P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}\otimes r_{h}^{\widetilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})\middle\|(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes r_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})\right)\right]\right].$ (C.4)
@@
C.2 Posterior Consistency
Lemma C.3.